OpenStudy (anonymous):

How do you solve a perfect square trinomial if it has a number in front of the first x value? For example: 25x^2+10x+1=9

OpenStudy (anonymous):

I know how to solve it if it was just x^2+10x+1=9, but i dont know what to do with the first 25

OpenStudy (anonymous):

5x^2+10x+1=9 5x^2+10x +1 - 9 = 0 5x^2+10x - 8 =0 you'll get a = 5, b = 10, and c = -8 substitute all of them to quadratic formula \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

If I put it into 5x^2+10x - 8 =0 that form wouldn't that be considered completing the square and i could just put all the x values on one side and divide each side by 25? then i could put it into b/2^2 to solve for x

OpenStudy (anonymous):

also how did you get the 5x^2 instead of the 25? @gerryliyana

OpenStudy (anonymous):

sorry wrong typo!

OpenStudy (anonymous):

i mean 25

OpenStudy (anonymous):

25x^2+10x+1=9 25x^2+10x +1 - 9 = 0 25x^2+10x - 8 =0 you'll get a = 25, b = 10, and c = -8 substitute all of them to quadratic formula

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] \[x = \frac{ -(10) \pm \sqrt{10^2 - 4*25*(-8)}}{ 2*25 } = \frac{ -10 \pm \sqrt{100 - (-800)} }{ 50 } \] \[x=\frac{ -10 \pm \sqrt{900} }{ 50 } = \frac{ -10 \pm 30 }{ 25 }\]

OpenStudy (anonymous):

So, you have \[x = \frac{ -10 + 30 }{ 50 } = \frac{ 20 }{ 50 } =\frac{ 2 }{ 5 }\] and \[x=\frac{ -10-30 }{ 50 } = \frac{ -40 }{ 50 } = -\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

oh , I understand now thank you!!

OpenStudy (anonymous):

you are very wlcome :)